### v1

@@ 5.1 Overall Performance Overview
(a) S&P 100 stock universe (2023-2024)
(b) S&P 500 stock universe (2024)
Figure 7: Cumulative returns of MarketSenseAI buy monthly signals against the market.
The attribution analysis (Table 7) reveals additional insights. With a 77–78% win rate across implementations, the system demonstrates remarkable consistency in signal precision. The positive alpha (17.6–18.9%) and elevated beta (1.24–1.27) of the S&P 500 portfolios suggest MarketSenseAI successfully identifies high-beta stocks with idiosyncratic upside potential. Furthermore, the stable monthly signal generation—35.1 ±7.95 buy signals for S&P 100 and 144.8 ±30.8 for S&P 500—indicates systematic selection rather than concentrated bets.
5.2 Factor Analysis and Risk Decomposition
To elucidate the drivers of MarketSenseAI’s outperformance, we decompose portfolio returns (MS-Eq - S&P 100) using the Carhart four-factor [52] and Fama-French five-factor [53] models. Both models explain a substantial portion of return variance ($R^{2}=88.4\%$ and $85.4\%$, respectively), validating their applicability. Key findings are summarized in Table 8 and discussed below.
5.2.1 Market Exposure and Size Bias
[…]
5.2.2 Value and Momentum Factors
Both models confirm consistent value exposure (HML = 0.08–0.11, $p<0.01$) [^7] High Minus Low (HML) is a value premium. It represents the spread in returns between companies with a high book-to-market value ratio (value companies) and companies with a low book-to-market value ratio., underscoring the Fundamentals Agent’s ability to identify undervalued equities through financial statement analysis. The Carhart model’s strong momentum loading (Mom = 0.18, p ¡ 0.01) highlights MarketSenseAI’s integration of price trends via the Dynamics Agent, a feature often absent in traditional fundamental models. This synergy between value and momentum aligns with the system’s architecture, where LLM-driven news sentiment and price dynamics reinforce fundamental insights.
5.2.3 Profitability and Investment Factors
The five-factor model reveals insignificant loadings on profitability (RMW) and investment (CMA) [^8] RMW (Robust Minus Weak) captures the profitability premium, while CMA (Conservative Minus Aggressive) represents the investment premium. These factors measure the spread in returns between companies with high versus low profitability, and conservative versus aggressive investment policies, respectively., suggesting these factors play minimal roles in MarketSenseAI’s strategy. This suggests MarketSenseAI’s returns are not systematically driven by these traditional style factors. The system’s integration of multiple data sources may help identify alpha sources beyond conventional factor premiums.
5.2.4 Alpha Generation and Unexplained Returns
[…]
Table 8: Factor Model Results
@@ 5.2.4 Alpha Generation and Unexplained Returns
[…]
88footnotetext: Note: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$. Dashes (–) indicate factor not included in model. Coefficients rounded to 3 decimal places. Market factor (Mkt-RF) shows near-unity exposure, SMB reflects large-cap bias, HML demonstrates value exposure, and Mom captures momentum effects.
6 Conclusion
@@
Future development will focus on two key directions: technological advancement through integration of reasoning-enabled LLMs (e.g., DeepSeek-R1, OpenAI’s O1) and market expansion to global and small-cap indices. These enhancements aim to further improve the system’s analytical capabilities while testing its adaptability across diverse market conditions.
MarketSenseAI represents a significant step forward in applying LLMs to financial analysis, offering both institutional and retail investors a transparent, data-driven approach to investment decision-making. By successfully addressing fundamental challenges such as processing lengthy documents, mitigating hallucination risks, and integrating multiple data sources, this work establishes a foundation for building more sophisticated, explainable investment frameworks.
Declarations
Acknowledgements
We acknowledge the financial support provided by the European Union Commission through Projects FAME [GA 101092639] and XR5.0 [GA 101135209], which contributed to the research presented in this paper.
Competing interests
The authors declare their involvement in the development and commercialization of the presented framework (i.e., MarketSenseAI) as a product of Alpha Tensor Technologies Ltd [^9] https://www.alpha-tensor.ai/.
Author contribution
George Fatouros: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing – original draft & Visualization. Kostas Metaxas: Conceptualization, Methodology, Software, Validation, Investigation, Data curation, Writing - Review & Editing. John Soldatos: Funding acquisition, Writing – review & editing. Manos Karathanasis: Resources & Supervision.

### v2

@@ V-B4 Alpha Generation and Unexplained Returns
TABLE VII: Factor Model Results
@@
Note: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$. Dashes (–)    indicate
factor not included in model. Market factor (Mkt-RF) shows    near-
unity exposure, SMB reflects large-cap bias, HML    demonstrates
value exposure, and Mom captures momentum effects.
VI Conclusions
@@
Future development will focus on two key directions: technological advancement through integration of reasoning-enabled LLMs and market expansion to global and small-cap indices. These enhancements aim to further improve the system’s analytical capabilities while testing its adaptability across diverse market conditions.
MarketSenseAI represents a significant step forward in applying LLMs to financial analysis, offering both institutional and retail investors a transparent, data-driven approach to investment decision-making. By successfully addressing fundamental challenges such as processing lengthy documents, mitigating hallucination risks, and integrating multiple data sources, this work establishes a foundation for building more intelligent investment frameworks.